Please help.


Which of the ordered pairs is a solution for the equation 2x-4y=8
(A) (-4,-2). (B) (0,-2), (C) (-4,0) or (D) (0,2)
Thanks
JMS

Answer by checkley71(8403)

answer is c) (-4,0)
proof:
2*-4-4*0=-8
-8=-8